# Cross-Camera Distracted Driver Classification through Feature Disentanglement and Contrastive Learning

### v1

@@ V Conclusions
Driver Monitoring Systems (DMSs) need to be robust to changes in vehicle configuration and camera positions. In this paper, we introduced DBMNet, a novel network designed for distracted driver classification, ensuring consistent performance regardless of camera placement. Given that DMSs operate on embedded devices with constrained computational resources, our model exploits GhostNet, a lightweight CNN, as the backbone. We improved the invariance to the camera view of the representation using a feature disentanglement module. In addition, we employed a supervised contrastive learning approach to increase consistency between intra-class representations and differentiate inter-class representations. Experimental results on the 100-Driver dataset demonstrated the effectiveness of the proposed model with an increment on average of 9% in Top-1 accuracy with respect to state-of-the-art methods. Through the analysis of features both before and after the feature disentanglement module, we determined that the objective of discarding view information is accomplished.
Some limitations are worth noting. Given the inherent complexity of the task, the accuracy results obtained are modest, indicating a need for further refinement to enhance the generalization capacity of our DBMNet. This requirement is particularly pronounced due to the low accuracy observed in the “Normal driving” class, which often gets confused with other distraction classes. Additionally, the inference time on our computational setup stands at 1.15 × $10^{-2}$ seconds, which remains relatively high for ensuring timely detection of driver distraction, hindering the implementation of effective safety systems.
Considering these identified limitations, as part of future endeavors, we aim to explore lighter backbones beyond GhostNet utilized in our experiments. This exploration aims to minimize the inference time while maintaining performance. Furthermore, we plan to investigate few-shot learning methods to calibrate the model to novel viewpoints, leveraging a limited number of training samples. These strategies are pivotal for addressing the current constraints and advancing the effectiveness of DMSs.

### v2

@@ V Conclusions
Driver Monitoring Systems (DMSs) need to be robust to changes in vehicle configuration and camera positions. In this paper, we introduced DBMNet, a novel network designed for distracted driver classification, ensuring consistent performance regardless of camera placement. Given that DMSs operate on embedded devices with constrained computational resources, our model exploits GhostNet, a lightweight CNN, as the backbone. We improved the invariance to the camera view of the representation using a feature disentanglement module. In addition, we employed a supervised contrastive learning approach to increase consistency between intra-class representations and differentiate inter-class representations.
Extensive experiments on the 100-Driver dataset demonstrated that DBMNet achieves a 9% improvement in Top-1 accuracy over state-of-the-art methods. Furthermore, cross-dataset evaluations on three benchmark datasets – AUCDD-V1, EZZ2021 and SFD – confirmed the strong generalization capabilities of our approach. Overall, the proposed method achieves a 7% gain in Top-1 accuracy compared to existing approaches. Through the analysis of features both before and after the feature disentanglement module, we determined that the objective of discarding view information is accomplished.
To validate real-world applicability, we deployed DBMNet and competing methods on a Coral Dev Board board. In this embedded setup, DBMNet demonstrated the lowest average error, while maintaining a compact model size, low memory footprint, fast inference time, and minimal power consumption.
Despite these strengths, some limitations remain. Due to the complexity of the task, overall accuracy is still modest, especially for “Normal driving” class, which is frequently misclassified as a distraction category. Additionally, the inference time of 1.15 × $10^{-2}$ seconds on board is relatively high for real-time safety-critical applications.
To address these issues, future work will focus on exploring lighter backbones beyond GhostNet to reduce inference time without sacrificing performance. We also plan to investigate few-shot learning strategies for model calibration on unseen camera viewpoints using limited labeled data. These improvements aim to enhance both the efficieny and adaptability of DBMNet in real-world DMS deployments.
Overall, our work highlights the critical importance of cross-camera generalization for real-world driver monitoring systems and provides a novel solution through a lightweight, disentanglement-based model rigorously evaluated under a realistic leave-one-camera-out protocol. This contributes a new perspective to the field by explicitly addressing viewpoint variability—an underexplored but practically crucial challenge.